OpenStudy (anonymous):

Write the equation of the line that passes through the point. (1,3);m=3

OpenStudy (anonymous):

\[y=m(x-x_{1})+y _{1}\] m=slope y=3(x-1)+3 y=3x-3+3 y=3x

OpenStudy (anonymous):

THNKS